### v1

# Reinforcing Trustworthiness in Multimodal Emotional Support Systems

Abstract
In today’s world, emotional support is increasingly essential, yet it remains challenging for both those seeking help and those offering it. Multimodal approaches to emotional support show great promise by integrating diverse data sources to provide empathetic, contextually relevant responses, fostering more effective interactions. However, current methods have notable limitations, often relying solely on text or converting other data types into text, or providing emotion recognition only, thus overlooking the full potential of multimodal inputs. Moreover, many studies prioritize response generation without accurately identifying critical emotional support elements or ensuring the reliability of outputs. To overcome these issues, we introduce MultiMood, a new framework that (i) leverages multimodal embeddings from video, audio, and text to predict emotional components and to produce responses responses aligned with professional therapeutic standards. To improve trustworthiness, we (ii) incorporate novel psychological criteria and apply Reinforcement Learning (RL) to optimize large language models (LLMs) for consistent adherence to these standards. We also (iii) analyze several advanced LLMs to assess their multimodal emotional support capabilities. Experimental results show that MultiMood achieves state-of-the-art on MESC and DFEW datasets while RL-driven trustworthiness improvements are validated through human and LLM evaluations, demonstrating its superior capability in applying a multimodal framework in this domain.
1 Introduction
Mental health challenges are an urgent global concern, profoundly affecting individuals and communities. The World Health Organization estimated that 970 million people—one in eight globally—lived with a mental disorder in 2019, primarily anxiety and depression (WHO). In the United States, nearly 23% of adults experienced mental illness in 2021 (NAMI). These conditions also impose a major economic burden, projected to reach \$6 trillion annually by 2030 (6billions). These figures underscore the need for scalable, innovative tools to support psychological well-being, with artificial intelligence (AI) emerging as a promising aid.
Figure 1: Example conversation illustrating the difference between prior systems and MultiMood. Prior methods respond with factual queries, whereas MultiMood demonstrates emotional awareness and offers empathetic, supportive feedback.
Advances in large language models (LLMs) and vision–language models (VLMs) have transformed text generation, dialogue systems, and multimodal reasoning (NEURIPS2020_1457c0d6; mitsui2024pslm). These models now support applications in healthcare (aihealthcare; li2023llava; nguyen2024logra), summarization (summarizing), and retrieval (airetrieval; airetrieval2). Their versatility makes them promising tools for addressing complex social challenges, including mental health care (malgaroli2025large; obradovich2024opportunities). Systems such as Woebot show that AI-driven dialogue can alleviate depression and anxiety through cognitive behavioral therapy (CBT)–inspired conversations (woebot; textonly1; textonly2). However, most existing systems remain text-only, overlooking nonverbal cues—tone, facial expression, and gesture—that are essential for empathy and trust (multiemotion1). Empirical studies indicate that multimodal signals strengthen emotional understanding and engagement in human–computer interaction (multiemotion2; saffaryazdi2025empathetic; loveys2020effect; eyben2015geneva; wang2022systematic). Consequently, text-only systems often lack the authenticity and nuance required for effective emotional support.
To address these limitations, we introduce MultiMood, a multimodal framework that integrates text, audio, and visual information to enhance emotional understanding in support-oriented dialogue (Figure 1). Unlike prior text-focused approaches, MultiMood leverages fine-grained cues—tone, prosody, facial expressions, and dialogue context—to generate empathetic, context-aware responses. Beyond multimodal fusion, MultiMood emphasizes trustworthy alignment: it employs reinforcement learning with human-defined psychological objectives to guide emotionally appropriate behavior. Specifically, it combines Proximal Policy Optimization (PPO) for stable learning with Group Relative Policy Optimization (GRPO) for fine-grained alignment to therapeutic standards, enabling fluent, safe, and ethically consistent responses suitable for AI-assisted emotional support.
MultiMood processes multimodal tokens to infer key emotional-support components—including user and supporter emotions, counseling strategies, and dialogue intent—which then guide response generation aligned with professional psychological frameworks. Our main contributions are as follows:
  • (i)
    Propose the MultiMood architecture, integrating multimodal features (text, audio, and vision) for emotional-support dialogue.
  • (ii)
    Design a trustworthiness-alignment framework with reinforcement-learning objectives that promote emotionally appropriate and reliable responses.
  • (iii)
    Evaluate state-of-the-art LLMs on a multimodal emotional-support dataset, demonstrating improvements in empathy, trustworthiness, and contextual accuracy.
Overall, this work advances the development of responsible multimodal emotional-support systems, offering a more holistic and human-centered approach to promoting psychological well-being.
2 Background
2.1 Dataset
The MESC dataset (smes), sourced from seasons 1-3 of In Treatment, comprises 1,019 dialogues and 28,762 utterances across text, audio, and video, annotated with 7 emotions (e.g., anger, sadness, disgust) and 10 therapeutic strategies (e.g., open questions, interpretation). Initially labeled using GPT-3.5 and refined by experts, it supports tasks like emotion recognition, strategy prediction, and response generation. MESC distinguishes itself from MELD (MELD) and ESConv (ESConv) with its multimodal and therapeutic focus; it advances empathetic AI for mental health. Analysis reveals prevalent neutral therapist emotions, reflecting their neutral stance to build client trust, which is consistent with counseling practices and not affecting model outcomes (smes). Besides, we also do experiments on the DFEW dataset (DFEW), a dynamic facial expression database. DFEW consists of over 16,000 video clips from movies, which were also annotated with seven emotions. These video clips contain various challenging interferences in practical scenarios such as extreme illumination, occlusions, and capricious pose changes.
2.2 Task Definition
Our goal is to emulate a human therapist’s nuanced functions in real-life therapeutic sessions. We decompose the AI-user interaction into four key tasks (smes) forming an emotionally intelligent support framework. Only Task 1 is referenced in both MESC and DFEW datasets, while Tasks 2–4 are exclusive to MESC:
  • (i)
    User Emotion Recognition (Task 1): Identifies the client’s emotion using multimodal cues (facial expressions, vocal prosody, text), enabling sensitive responses to psychological needs.
  • (ii)
    System Emotion Prediction (Task 2): Predicts the system’s emotional tone (e.g., neutral, angry,…) to align with the chosen strategy, fostering rapport and trust.
  • (iii)
    System Strategy Prediction (Task 3): Selects the optimal therapeutic strategy (e.g., validation, reflection) based on user emotion and dialogue history, mirroring tailored therapist techniques.
  • (iv)
    System Response Generation (Task 4): Generates a natural, contextually appropriate response embodying the predicted tone and strategy, promoting emotional safety and insight.
2.3 Related works
Framework Approach Training method Output
Visual Audio Text User Emo. Therapist Emo. Strategy Response Trust. Aware.
InternVideo2.5 SFT+RL
VideoLLaVA SFT
EmotionLLaMA SFT
SMES SFT
MultiMood (ours) SFT+RL
Table 1: Comparison between MultiMood and other multi-LLM models for emotion recognition. “SFT" = supervised fine-tuning, “RL" = reinforcement learning, “Resp." = response generation, “Trust" = trust-awareness. InternVideo2.5 has never been used for emotional tasks before.
Emotional Support Frameworks
In psychological counseling, several established theoretical frameworks guide practitioners in addressing psychological and emotional difficulties. CBT (CBT) is a structured, evidence-based approach that targets maladaptive thoughts to improve emotions and behaviors, ideal for anxiety and depression. Psychodynamic Therapy (psychodynamic) explores unconscious conflicts and early experiences to enhance self-awareness, suited for issues like personality disorders. Humanistic Therapy, such as Rogers’ person-centered (humanistic) approach, fosters self-actualization through empathy and unconditional regard, effective for self-esteem and existential concerns. Acceptance and Commitment Therapy (ACT) (ACT), a mindfulness-based cognitive approach, promotes psychological flexibility by encouraging acceptance and value-driven actions, applicable to conditions like chronic pain and anxiety.
Recent advancements in AI have enhanced emotional support systems, addressing limitations of smartphone-based conversational agents (smartphone; milne2022artificial), Muffin framework (muffin) uses model-agnostic AI feedback and contrastive learning to improve response fluency and relevance. Hybrid Empathetic Framework (HEF) (HEF) integrates LLMs with small-scale empathetic models to enhance emotion detection and response generation. The Sequential SMES framework (smes) leverages multimodal data to simulate therapeutic empathy and deliver tailored responses. Our MultiMood framework advances these efforts by incorporating trustworthiness through reinforcement learning with PPO and GRPO, ensuring safe, empathetic, and contextually appropriate responses for diverse user needs.
Multimodal LLMs
Multimodal LLMs like InternVideo2.5 (internvideo2.5) and VideoLLaVA (videollava) advance video understanding. InternVideo2.5 employs a single InternVIT encoder with Hierarchical Token Compression (HICO) to efficiently process long videos, merging similar tokens to reduce computation while preserving quality (internvideo2.5). Its three-stage training supports tasks like temporal grounding and object tracking. VideoLLaVA aligns images and videos using LanguageBind for unified visual representation via a shared projection layer (videollava), but lacks audio processing, unlike MultiMood’s modality-specific projectors. These models focus on visual content while missing audio cues (volume, tone, pitch) critical for emotion recognition. Multimodal LLMs also enhance emotional support, overcoming single-modality limitations by capturing nuanced emotional signals for empathetic AI. EmotionLLaMA (emotionllama) uses the MERR dataset (28,618 samples) and specialized encoders for precise emotion recognition, surpassing GPT-4V in zero-shot tasks (zhang-etal-2024-visual). The SMES framework, with the MESC dataset (28,762 utterances from In Treatment), processes multimodal inputs for emotion recognition, strategy prediction, and response generation, improving therapeutic mimicry (smes). MultiMood stands out with specialized encoders per modality and a reinforcement learning algorithm designed for trustworthiness, as summarized in Table 1.
Trustworthiness in Responses
Trustworthiness is essential for effective emotional support from therapists and doctors, fostering a safe space for patient vulnerability. Goleman’s emotional intelligence framework (goleman) emphasizes empathy, self-regulation, and social skills as key to building trust, enabling clinicians to communicate effectively. Crits-Christoph et al. (client-clinician) highlight that trust, distinct from therapeutic alliance, encourages sharing private information, with racial disparities (e.g., lower trust among Black patients) underscoring equity’s role. Richmond et al. (trustindoctor) link trustworthiness to communication, fidelity, and fairness, noting that lower trust can delay care. In LLMs, trustworthiness is critical for safe, supportive interactions, as outlined in TrustLLM’s eight dimensions (trustllm): truthfulness ensures accuracy, safety fosters healthy dialogue, fairness promotes impartiality, and robustness ensures reliability. Privacy protects autonomy, machine ethics ensures moral behavior, transparency provides clarity, and accountability holds LLMs responsible. In MultiMood, these factors are integrated to train robust LLMs, significantly reducing hallucination. Building on these foundations, we propose a tailored set of trustworthiness dimensions for emotional support systems to improve automatically generated responses to meet therapeutic standards.
3 Methodology
3.1 Overview
Figure 2: MultiMood overview. Multimodal architecture that processes video, audio, text, and historical conversation data through dedicated encoders. The modality-specific embeddings are fused and passed into an LLM, which is further optimized using reinforcement learning guided by trustworthiness criteria to generate emotionally supportive and responsible responses.
The MultiMood framework, shown in Figure 2, integrates an audio encoder $\mathcal{E}^{aud}$, a visual encoder $\mathcal{E}^{vis}$, a conversation compressor $\mathcal{C}$, and a large language model $\phi$. For an input tuple $P=\langle\text{Audio},\text{Video},\text{Prompt},\text{History}\rangle$, the model is defined as:
$\hat{O}=\Psi(\phi,\mathcal{E},\Omega,\mathcal{C},P),$ (1)
where $\mathcal{E}$ combines audio, vision, and text encoders, $\Omega$ is the vision pre-processor, and $\hat{O}$ is the text output. A multi-tower architecture generates modality-specific embeddings: video via vision tower $f_{V}$, audio via audio tower $f_{A}$, and text via $\phi$ tokenizer $f_{T}$, yielding $E_{T}=f_{T}(\text{Prompt})$. A compressor distills text histories into concise representations, $E_{H}=\text{ConvCompressor}(H)$, enabling efficient context processing. Embeddings $E^{\prime}_{V};E^{\prime}_{A};E_{T};E^{\prime}_{H}]$ are aligned via modality-specific projectors and fed into the LLM to predict outcomes for four tasks.
3.2 Framework Components
Modality-Specific Encoder
The vision pre-processor $\Omega$ uses the input video as a frame sequence, processed by a CLIP-based (CLIP) visual encoder $\mathcal{E}^{vis}$ to extract video features:
$E_{V}=\mathcal{E}^{vis}(\Omega(\text{Video})).$ (2)
A Spatial-Temporal Convolution (STC) connector (videollama2) captures spatial and temporal dynamics:
$E^{\prime}_{V}=\text{STC}(E_{V})=P_{V}(\text{R}_{2}(\text{Conv3D}(\text{R}_{1}(E_{V})))),$ (3)
where $\text{STC}(\cdot)$ includes two spatial interaction modules ($\text{R}_{1}$, $\text{R}_{2}$) and a 3D convolution (Conv3D), with $P_{V}$ projecting features to the language model $\phi$ space.
For audio, the BEATs model (beats) serves as the audio encoder $\mathcal{E}^{aud}$, extracting features mapped to the language model space via a linear projector $P_{A}$:
$E_{A}=\mathcal{E}^{aud}(\text{Audio}),\quad E^{\prime}_{A}=P_{A}(E_{A}).$ (4)
Cross-Modality Concatenation
Our approach draws from ECoT (ECoT), a plug-and-play method that boosts LLM performance in emotional generation tasks by aligning with Goleman’s emotional intelligence theory (goleman1995emotional). We design a prompt template to guide the LLM in generating empathetic responses, integrating historical and real-time data. Multimodal features are concatenated into the input using specialized tokens: <video>, <audio>, and <history>, replaced by processed embeddings $E^{\prime}_{V}$, $E^{\prime}_{A}$, and $E^{\prime}_{H}$, respectively, forming the input sequence $X_{LLM}$. This attention-based fusion enables the model to dynamically prioritize relevant cues (e.g., tone, facial expressions) for safe, context-aware responses, while simultaneously predicting three classifications and generating therapist-like outputs (see Figure 2). More specific concatenation is discussed in the Appendix A.1.
Conversation Compressor
Figure 3: ConvCompressor architecture and pretraining.
Conversational emotional support systems require effective processing of extensive dialogue histories to deliver contextually appropriate responses. However, long conversation histories pose computational and memory challenges for language models. To address this, we propose Conversation Compressor (ConvCompressor), a lightweight module that distills dialogue histories into compact, semantically rich representations while retaining critical information. ConvCompressor employs the Mamba state-space model (mamba) as its core, offering linear computational complexity compared to the quadratic scaling of transformers. It appends a <MEM> token to each conversational turn $U_{i}$ in a history $H=\text{concat}(\{U_{i}\}_{i=1}^{T})$, where $U_{i}$ includes role information, utterance content, emotional labels, and therapist strategy labels, forming $H^{\prime}=U_{1}\texttt{<MEM>}U_{2}\texttt{<MEM>}\ldots U_{T}\texttt{<MEM>}$. The Mamba backbone processes $H^{\prime}$ to generate hidden representations $Z$, from which we extract hidden states only at <MEM> token positions. The extracted representations then undergo a trainable memory projector $P_{H}$ before being fed to LLM (see Figure 3).
ConvCompressor is optimized through a two-stage training process. First, it is pre-trained with a frozen language model on a reconstruction task to regenerate the original dialogue history from compressed <MEM> representations. Then, it undergoes end-to-end fine-tuning within the multimodal pipeline, adapting its compression strategy to significantly reduce the number of input tokens for the LLM while preserving a comparative overall performance.
Trustworthiness Dimensions for Emotional Support
Dimension Source Definition Truthfulness (trustllm) (trustindoctor) (goleman) The accurate representation of information, facts, and results by the AI system. Safety (trustllm) (goleman) Promote safe, healthy conversations, avoiding harm, distress, or triggers while supporting user well-being. Fairness (trustllm) (trustindoctor) (goleman) The quality of being impartial and equitable, considering multiple perspectives and maintaining a positive, action-oriented tone. Privacy (trustllm) (trustindoctor) Practices that safeguard human autonomy, identity, and data dignity. Empathy (trustindoctor) (goleman) Openness and honesty in expressing sympathy for negative situations or approval for positive ones. Reliability (client-clinician) (trustindoctor) (goleman) Responses foster understanding, connection, and provide encouragement, comfort, or support. Ethical Guidance (trustllm) (goleman) Ensuring AI behaviors guide emotional health responsibly, avoiding manipulation or harm.
Inter-annotator Agreement (Top)
Dimension Flu. Ide. Com. Sug. Ove.
Fleiss Kappa 0.65 0.61 0.60 0.61 0.67
Human Evaluation (Middle)
Flu. Ide. Com. Sug. Ove.
Qwen2-7B 22% 17% 17% 22% 21%
MM (SFT) 23% 30% 27% 25% 21%
MM (SFT+RL) 55% 53% 56% 53% 58%
LLMs Evaluation (Bottom)
Model Judge: GPT-4o
Tru. Saf. Fai. Pri. Emp. Rel. Eth. Avg.
Qwen2-7B 6.0 4.2 5.1 8.0 4.2 4.6 4.3 5.2
MM (SFT) 6.2 4.3 5.8 7.8 4.3 4.8 4.9 5.4
MM (SFT+RL) 7.0 6.3 6.8 8.8 6.2 6.4 6.3 6.8
Judge: Claude 4.0-Sonnet
Qwen2-7B 4.9 7.0 6.0 8.0 5.6 7.0 6.0 6.4
MM (SFT) 5.0 7.0 5.9 8.0 5.9 7.3 6.0 6.5
MM (SFT+RL) 7.2 7.0 6.5 8.0 7.8 7.4 6.6 7.2
Judge: Grok-3
Qwen2-7B 6.1 6.0 7.2 7.5 6.0 6.3 5.8 6.4
MM (SFT) 6.2 6.6 7.6 7.2 6.1 6.9 5.9 6.6
MM (SFT+RL) 7.3 7.8 8.7 9.3 7.5 7.9 7.5 8.0
Table 2: Trustworthiness dimensions for emotional support tasks (left); Inter-annotator agreement, human and evaluation results (right) across different models. Flu., Ide., Com., Sug., Ove., stand for Fluency, Identification, Comfort, Suggestions, and Overall, whose definitions are provided in the Appendix C.1.
3.3 Training
The training process comprises two key stages that enhance both robustness and trustworthiness.
Stage 1: Supervised Fine-Tuning We fine-tune our framework on the MESC dataset, leveraging multimodal data throughout the training process. To accommodate potential missing modalities (video or audio) during inference, we introduce a random modal selection mechanism. This is defined by a probability vector $\mathbf{p}=[p_{a},p_{v},p_{av}]$, representing the likelihoods of selecting audio or video or both modalities. This approach enhances the framework’s robustness by exposing it to all possible modality combinations during training. For multimodal processing, we employed SigLIP-So400M-Patch14 384 (DBLP:siglip/iccv/ZhaiM0B23) for video, and BEATs (beats) for audio. The ConvCompressor is built on Mamba-370M (mamba). Prompt for training is provided in the Appendix A.1.
Stage 2: Trustworthiness-Aware via Reinforcement Learning Initial assessments showed that the system’s responses occasionally lacked the natural flow and trustworthiness needed for effective emotional support. To address this, we defined a set of trustworthiness criteria, $\mathcal{C}=\{c_{1},c_{2},\dots,c_{k}\}$ (see Table 2-left), and employed reinforcement learning to align responses with ethical and therapeutic standards. We used Group Relative Policy Optimization (GRPO) (grpo) and Proximal Policy Optimization (PPO) (ppo) after supervised fine-tuning to enhance response quality. GRPO optimizes by comparing responses within groups, while PPO stabilizes learning through clipped updates and a Kullback-Leibler divergence penalty. To guide learning, we designed a reward function combining trustworthiness and similarity. The similarity score $r_{\text{sim}}(y,y^{*})$ leverages BGE-M3 embeddings via ColBERT (Colbert), integrating dense, sparse, and ColBERT-specific similarities with weights (1, 0.3, 1), normalized to $[0,1]$:
$r_{\text{sim}}=\text{scale}\left(\text{sim}_{\text{C}}+0.3\,\text{sim}_{\text{s}}+\text{sim}_{\text{d}}\right)$ (5)
Trustworthiness $r_{\text{trust}}(y)$ is evaluated by GPT-4o per sentence, averaged and scaled to $[0,1]$. GPT-4o is trusted for this task due to its proven capabilities in labeling data (llm-anno) across various tasks, as well as its robust safety mechanisms  (safety-mechanisms):
$r_{\text{trust}}(y)=\text{scale}\left(\frac{1}{|y|}\sum_{i=1}^{|y|}\text{GPT-4o}_{\text{trust}}(y_{i})\right)$ (6)
The final reward is:
$r(y|x)=\frac{1}{2}\left(r_{\text{trust}}(y)+r_{\text{sim}}(y,y^{*})\right)$ (7)
Details of RL strategies are in the Appendix A.2.
Method Hap Sad Neu Ang Sur Dis Fea UAR WAR
IAL (IAL) 87.95 67.21 70.10 76.06 62.22 0.00 26.44 55.71 69.24
VideoMAE (videoMae) 93.09 78.78 71.75 78.74 33.44 17.93 41.46 63.60 74.60
S2D (S2D) 93.62 80.25 77.14 81.09 64.53 1.38 34.71 61.82 76.03
EmotionLLaMA (emotionllama) 93.05 79.42 72.47 84.14 72.79 3.45 44.20 64.21 77.06
MultiMood (ours) 96.31 93.68 89.45 88.82 81.68 78.38 85.19 85.94 89.89
Table 3: Comparison of multimodal emotion recognition results on DFEW.
3.4 Trustworthiness Dimension Table
To assess response trustworthiness in emotional support, we first developed a domain-specific framework, Trustworthy Dimensions. This was built by synthesizing insights from four key sources: the TrustLLM framework (trustllm), which outlines trust principles for LLMs; patient-clinician trust studies (client-clinician; trustindoctor); and Goleman’s emotional intelligence principles (goleman). From TrustLLM, we adopted core technical values such as Truthfulness, Safety, Fairness, Privacy, and Machine Ethics. Clinical trust literature contributed Honesty, Communication, Confidentiality, Fidelity, and Reliability - emphasizing relational trust. Goleman’s work added Empathy and Social Skills, highlighting emotional resonance. These elements were distilled into seven core dimensions, carefully defined to balance technical reliability with emotional sensitivity (Table 2-left). Details of each criterion, informed by prior literature, are provided in the Appendix A.3.
4 Experiments
4.1 Experimental Setup
Metrics
For generation evaluation, we use BLEU-n (B-2), ROUGE-L (R-L), and BERTScore (BS) to evaluate the Therapist’s responses from the models. For classification of the MESC dataset (smes), we use Accuracy and Weighted-F1 as metrics. These metrics collectively provide a comprehensive overview of model performance across different tasks. For the DFEW dataset (DFEW), we use unweighted average recall (UAR) and weighted average recall (WAR) to compare our method with SOTA methods.
Baselines
We utilized the pretrained LLM from VideoLLaMA2 (videollama2) as the multimodal LLM backbone, leveraging its training on multimodal data. We compare MultiMood with API-based LLMs (GPT-4o (GPT4), Grok3 (xai2025grok2), Claude-3.7 (claude), Deepseek-R1 (deepseek), LLaMA4 (llama4)); Open-source LLMs (Qwen2 (qwen2) and Qwen2.5 (qwen25)) - finetuned on the MESC dataset (smes) with SFT and PPO; Multimodal frameworks (EmotionLLaMA (emotionllama), VideoLLaMA2, VideoLLaVA (videollava), InternVideo2.5-8B (internvideo2.5), VideoLLaMa3-7B (videollama3)) - not finetuned due to resource constraints; and SMES-leveraged models (MMGCN (MMGCN), MMDFN (MMDFN), Blenderbot (ESConv)).
Settings
[…]
4.2 Results
In this section, we present a comprehensive evaluation to compare our framework with other frontier models on the MESC and DFEW datasets. The evaluation highlights the strengths and advancements of our framework in handling complex multimodal data in both tasks. Ablation study of modalities affect show in Appendix D.
Model Training method Modality Task 1 Task 2 Task 3 Task 4
Acc F1 Acc F1 Acc F1 B2 R-L BScore
MMGCN SFT A,V,T 55.80 57.58 - - - - - - -
MMDFN SFT A,V,T 58.13 55.86 - - - - - - -
Blenderbot SFT SFT A,V,T - - - - 48.00 46.10 1.31 15.38 86.60
SMES SFT A,V,T 54.60 46.80 96.10 64.00 49.00 20.20 5.13 15.42 86.80
VideoLLaMA2-72B - A,V,T 55.06 55.68 97.36 98.10 25.77 26.09 3.55 13.77 85.37
VideoLLaVA SFT V,T 46.60 47.08 94.18 88.03 27.31 22.28 4.37 9.84 84.23
InternVideo2.5-8B SFT V,T 37.22 34.69 98.90 98.79 13.44 4.82 3.92 13.21 85.40
VideoLLaMA3-7B SFT A,V,T 45.28 46.23 97.40 72.66 33.96 24.50 3.70 11.55 85.07
EmotionLLaMA SFT A,V,T 46.12 41.95 99.11 99.11 37.44 25.41 2.55 10.76 84.28
LLaMA4-Maverick - T 23.34 21.16 68.72 81.02 14.53 8.11 3.94 10.03 84.26
Claude-3.7-Sonnet - T 32.59 33.33 85.90 91.80 27.97 27.55 2.25 8.45 83.79
Deepseek-R1 - T 20.48 20.27 59.47 74.03 17.84 15.95 3.22 9.20 83.96
GPT-4o - T 38.98 43.56 72.46 83.60 24.88 26.26 2.30 9.20 84.31
Grok-2 - T 22.46 25.08 65.85 78.80 20.44 18.19 2.29 9.60 84.61
Qwen2-7B SFT A,V,T 41.83 37.16 99.33 99.00 37.43 33.52 4.68 13.20 85.61
Qwen2-0.5B SFT+PPO A,V,T 44.27 44.99 99.33 99.00 36.34 33.01 4.40 12.31 85.36
Qwen2-7B SFT+Comp. A,V,T 44.27 44.03 99.33 99.00 39.42 35.69 4.60 12.90 85.47
Qwen2.5-7B SFT A,V,T 53.00 51.13 98.63 98.80 35.14 34.46 4.68 13.81 85.52
MultiMood SFT+Comp. A,V,T 53.75 51.75 99.33 99.00 39.29 36.25 5.26 15.34 85.81
MultiMood SFT A,V,T 56.38 55.81 99.11 99.11 36.78 34.32 4.58 13.47 85.71
MultiMood SFT+GRPO A,V,T 58.60 57.78 99.33 99.00 42.81 39.65 6.18 17.86 86.80
MultiMood SFT+Comp+GRPO A,V,T 55.94 55.33 99.11 99.11 38.10 34.58 5.42 15.83 86.00
Table 4: Benchmark of MultiMood against other baselines on MESC. Task 1: User Emotion Recognition, Task 2: System Emotion Recognition, Task 3: Strategy Prediction, Task 4: Response Generation. A=Audio, V=Video, T=Text; B2=BLEU-2; R-L=ROUGE-L; BScore=BERTScore (F1); Comp.=Conversation Compressor.
Overall Performance
Tables 3 and 4 present the primary results of our proposed MultiMood framework compared to baseline models, evaluated across four MESC tasks (smes) and one DFEW task (DFEW). MultiMood demonstrates exceptional adaptability, achieving robust performance across all tasks, unlike baseline models that often excel in specific domains. It delivers consistent results in emotion recognition, strategy prediction, system emotion prediction, and response generation, surpassing larger models like VideoLLaMA2-72B and specialized classification models like MMGCN. Notably, MultiMood achieves the highest average score (56.45) across the four MESC tasks and a SOTA score on the DFEW dataset. The ConvCompressor module demonstrates remarkable efficiency, achieving 98.6% token reduction (see result in Appendix B.2) while maintaining competitive performance, making our framework significantly more memory-efficient for processing extended dialogue histories. Our framework performance is evaluated from four key perspectives.
Emotion Recognition: Our MultiMood framework achieves SOTA performance on the single-labeled DFEW dataset (DFEW), outperforming prior methods in accuracy, unweighted average recall and weighted average recall scores, as shown in Table 3. It achieves the highest UAR of 85.94% and WAR of 89.89%, excelling across all emotion categories, notably Disgust (78.38%), where prior models like IAL (IAL), VideoMAE (videoMae), S2D (S2D), and EmotionLLaMA (emotionllama) struggled due to under-representation (DFEW). With MESC, the variant utilizing GRPO attains the highest performance, followed closely by the fine-tuned framework without GRPO. MultiMood surpasses video understanding models (e.g., VideoLLaMA, InternVideo2.5), the Qwen family, and closed-source models, as well as specialized frameworks like SMES (smes) and MMDFN (MMDFN) (shown in Table 4). MultiMood’s robust classification, particularly for nuanced emotions, enhances empathetic response generation, establishing a new benchmark for precise emotion recognition. However, while ConvCompressor improves memory efficiency, it may compromise performance due to information loss.
Strategy Prediction: MultiMood achieves a 42.81% accuracy on the Strategy Prediction task, slightly trailing BlenderBot SFT (48%) and SMES (49%) (smes). This gap reflects MultiMood’s design prioritizing robust, generalized performance across diverse tasks over specialization in strategy prediction. Nonetheless, it delivers a competitive F1 score, surpassing several baselines, though marginally behind SMES in accuracy. Unlike BlenderBot, which benefits from domain-specific retrieval tools, MultiMood faces challenges with class imbalance. However, its instruction-guided framework excels in generating safe, multimodal-aware responses, enhancing generalizability across emotion recognition, strategy planning, and empathetic response generation.
System Emotion Prediction: Most fine-tuned models achieve over 90% accuracy in this task, attributed to a data skew where 90% of labels are Neutral. This imbalance is typical in emotional support contexts, as therapists maintain a calm demeanor, enables the system to generate honest, unbiased answers.
Response Generation: MultiMood (SFT+GRPO) achieves superior performance across all metrics–BLEU-2 (6.18), ROUGE-L (17.86), and BERTScore (86.80)–demonstrating the efficacy of combining Group Relative Policy Optimization with supervised fine-tuning to produce fluent, contextually aligned responses. It outperforms baselines like VideoLLaMA2-7B (SFT) and Qwen2-7B (SFT + PPO), as well as closed-source models such as GPT-4o (GPT4) and LLaMA4 (llama4), which underperform due to their reliance on textual features alone. MultiMood’s integration of multimodal data enhances its classification and response generation capabilities, surpassing recent SOTA SMES (smes) and setting a new benchmark for empathetic, high-quality responses. Some examples of response show in Appendix E.
Human and LLM Evaluation
We conducted a comprehensive evaluation using both human and LLM assessments to assess the trustworthiness and quality of responses from Qwen2-7B (SFT), MultiMood–MM(SFT), and MultiMood–MM(SFT+RL). Four graduate students served as human annotators, all with expertise in emotional support research and advanced English proficiency (IELTS overall $\geq$ 7.0 with reading $\geq$ 7.5) to ensure accurate evaluation of text-only outputs. They received training with tutorials and examples, including framework-generated outputs, dialogue contexts, situational details, and responses from a licensed psychologist, followed by a test on 100 MESC dataset validation samples to achieve a Cohen’s kappa inter-annotator agreement above 0.4 (cohenkappa) (see Table 2-right-middle); retraining was required if unmet. During annotation, two annotators labeled all responses, with discrepancies resolved by a third and persistent disagreements settled by a fourth to establish the majority label, detailed results in Table 2-right-middle. The annotation guideline is provided in the Appendix C.1.
Human evaluation shows MultiMood (SFT+GRPO) outperforming in Fluency (55%), Comfort (56%), and Overall (58%), highlighting the effectiveness of multimodal fine-tuning and GRPO in enhancing response quality. Simultaneously, LLM evaluation, guided by (llm-anno), (llm-medical-1), and (llm-medical-2), underscored LLMs’ near-human accuracy in surgical knowledge but noted query inconsistency, stressing stable evaluation needs. LLM scoring pre- and post-application of our trustworthiness dimension table 2-left revealed RL-incorporated frameworks significantly outperformed non-RL frameworks across three LLMs (see Table 2-right-bottom). By aligning with trustworthiness criteria, RL enhances safety, reliability, and ethical soundness, addressing non-RL inconsistencies and boosting utility for critical applications. Prompt for LLMs evaluation is provided in the Appendix C.2.
5 LIMITATION
Despite the demonstrated effectiveness of our framework, several limitations persist. It underperforms BlenderBot in strategy prediction (per SMES (smes)) due to class imbalance and the lack of external retrieval. The inability to fine-tune certain multimodal frameworks, constrained by resource limitations, weakens the robustness of our comparisons. Additionally, although its usage was proved (llm-anno), using GPT-4o for trustworthiness evaluation may introduce bias, particularly when its reward function influences reinforcement learning training. Furthermore, the experimental datasets, derived from movies and challenges rather than real treatment settings, lack authenticity—a common issue in this field (medical-data-1; medical-data-2), underscoring the need for more realistic emotion support datasets in future research.
6 CONCLUSION
In conclusion, MultiMood leverages multimodal techniques to achieve state-of-the-art results in emotion recognition and response generation, outperforming closed- and open-source models. Enhanced by reinforcement learning, it demonstrates high trustworthiness per human and LLM evaluations, with potential for therapeutic use. However, limitations in strategy prediction, hardware constraints and lack of realistic datasets suggest areas for future enhancement.
7 Acknowledgments
This research was supported by The Mohamed bin Zayed University of Artificial Intelligence’s Travel Grant and the VNUHCM-University of Information Technology’s Scientific Research Support Fund.
Appendix A Training
A.1 Training Prompt for SFT
To enable our framework to process information from diverse modalities, we incorporate specialized <video>, <audio>, and <history> tokens into the multimodal large language model input, representing video, audio, and historical embeddings, respectively. Specifically, the Vision tower $E_{V}^{\prime}$ and Audio tower $E_{A}^{\prime}$ substitute the <video> and <audio> tokens, while the ConvCompressor $E^{\prime}_{H}$ replaces the <history> token in the input text prompt, creating the input sequence $X_{LLM}$. The model is designed to process $X_{LLM}$ and simultaneously predict three classification outcomes while generating a therapist-like response, informed by insights from these classifications. We utilize the following sample template:
[CONTEXT]
Problem: {problem_type}
Situation: {situation}
[CURRENT CONTEXT]
Video: <video>
Audio: <audio>
Chat history: <history>
Client utterance: {user_utterance}
[PROMPT]
Within this structured input, our LLM dynamically prioritizes the most relevant information across modalities, adapting to the emotional context and task. This attention-based fusion allows the model to optimally combine cues like tone, facial expressions, and text semantics, ensuring safe, empathetic, and contextually appropriate responses.
As illustrated in Figure 4, our full prompt draws inspiration from ECoT (ECoT), a plug-and-play prompting technique that boosts LLM performance on emotional generation tasks by aligning with human emotional intelligence principles. These include Social Skills: Influencing others’ emotions, Self-Regulation: Controlling negative self-emotions, Self-Awareness: Recognizing self-emotions, Empathy: Recognizing others’ emotions, and Motivation: Activating positive self-emotions. We guide the LLMs to generate emotional responses in conversations with context, following an emotional thinking process based on Goleman’s Theory (goleman1995emotional).
{promptbox}
[CONTEXT]
Problem: problem_type
Situation: situation
History chat information above
[CURRENT CONTEXT]
Video: <video>
Audio: <audio>
Chat history: <history>
Client utterance: user_question
The [CONTEXT] is the history of the current conversation between ’Client’ and ’Therapist’. [CURRENT CONTEXT] is the current ’Client’ turn.
Now, as the ’Therapist’, predict the Client’s emotion, Therapist’s emotion, and strategy, then craft an empathy response. Follow these steps:
Step 1: Understand the context and conversation content.
Step 2: Predict and justify: - Client’s emotion: (anger, sadness, disgust, depression, neutral, joy, fear). - Therapist’s emotion: (anger, sadness, disgust, depression, neutral, joy, fear). - Therapist’s strategy: (open question, approval, self-disclosure, restatement, interpretation, advisement, communication skills, structuring the therapy, guiding the pace and depth, others). Guide: Communication Skills: Small talk and body language; Advisement: Guidance or solutions; Structuring: Set therapy goals; Guiding: Regulate conversation flow; Others: Other strategies.
Step 3: Craft an empathy response using Therapist’s emotion and strategy, aligning with Client’s perspective, avoiding negative triggers, and promoting well-being.
Step 4: Revise the response, avoid hurting feelings, and consider response impact.
[OUTPUT FORMAT]
Client’s emotion:
Therapist’s emotion:
Therapist’s strategy:
Therapist’s response:
Figure 4: Full prompt used for training and inference
A.2 Reinforcement Learning Strategies
We implement GRPO and PPO as our RL algorithms to supervise LLM models given trustworthy conditions, as in Table 5.
Group Relative Policy Optimization (GRPO) (grpo) is a reinforcement learning method tailored for training Large Language Models (LLMs) as policies. Instead of relying on a value-based critic model, GRPO computes relative advantages within a group of completions sampled from a prompt. For each question, the model generates a set of answers. The answers are then scored by the reward functions. Based on those scores the model avoids low-scoring answers and is encouraged to correct errors to generate high-scoring answers, thereby improving the inference ability. GRPO has achieved a significant improvement on math tasks. At the same time, not using the critic model helps to reduce a large amount of computational resources (deepseek).
In GRPO, For each question $q$, GRPO samples a group of outputs $\{o_{1},o_{2},\cdots,o_{G}\}$ from the old policy $\pi_{\theta_{\text{old}}}$ and then optimizes the policy model by maximizing the following objective:
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O|q)}\left[\mathcal{L}_{\text{GRPO}}(\theta)\right]$ (8)
$\mathcal{L}_{\text{GRPO}}(\theta)=-\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}l_{i,t}\\$ (9)
$l_{i,t}=\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta}(o_{i,t}|q,o_{i,<t})_{\text{no grad}}}\hat{A}_{i,t}-\beta D_{\text{KL}}[\pi_{\theta}||\pi_{\text{ref}}]\\$ (10)
For each of the $G$ sequences, GRPO compute the reward using a reward model. To align with the comparative nature of reward models–typically trained on datasets of comparisons between outputs for the same question–the advantage is calculated to reflect these relative comparisons. It is normalized as follows:
$\hat{A}_{i,t}=\frac{r_{i}-\text{mean}(r)}{\text{std}(r)}$ (11)
KL divergence is estimated using the approximator introduced by (net2020kl). The approximator is defined as follows:
$D_{\text{KL}}[\pi_{\theta}||\pi_{\text{ref}}]=\frac{\pi_{\text{ref}}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}-\log\frac{\pi_{\text{ref}}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}-1$ (12)
$\beta$} parameter to adjust the KL penalty to prevent the model from straying too far from the initial policy.
To guide the training of our GRPO algorithm for optimizing large language models, we employ a reward function that balances trustworthiness and semantic similarity, as detailed in Section 3.3. The similarity score $r_{\text{sim}}(y,y^{*})$ utilizes BGE-M3 embeddings via ColBERT (Colbert), combining dense, sparse, and ColBERT-specific similarities with weights (1, 0.3, 1), normalized to $[0,1]$. Trustworthiness $r_{\text{trust}}(y)$ is computed by averaging per-sentence GPT-4o trust scores, scaled to $[0,1]$. The final reward, given in Equation 7, averages these two components, as described above.
Proximal Policy Optimization (PPO) (ppo) is a widely adopted reinforcement learning algorithm for fine-tuning large language models via Reinforcement Learning from Human Feedback (RLHF). Similar to GRPO, PPO begins with a rollout phase where the model generates responses to input prompts, followed by evaluation using a reward model. It incorporates a KL-divergence penalty to constrain policy shifts and employs Generalized Advantage Estimation (GAE) (gae) to guide which token probabilities should be reinforced. A key feature of PPO is its clipping mechanism, which stabilizes learning by preventing drastic updates between old and new policies. For GRPO (grpo), the LLM generated four responses per prompt, evaluated using the reward function in Equation 7. GRPO optimizes based on both reward signals and the KL divergence between the current and reference models. We set the KL parameter to 1.12 to balance exploration and stability.
For PPO, we first trained a reward model on sentence pairs (<chosen>, <rejected>), where <chosen> represents ground truth references from the original dataset, and <rejected> consists of LLM-generated responses with ROUGE-L scores below 0.3 when compared to references. Using this model, PPO fine-tuned the LLM by updating both policy and value networks via advantage-based loss, with a clipping threshold of 0.2 to prevent unstable updates. Due to computational constraints, we were unable to apply PPO to our 3B model.
A.3 Creation of Trustworthiness Dimensions Table
Dimension Source of Derivation Definition
Truthfulness (trustllm), Honesty of (trustindoctor), Empathy of (goleman) Accurate representation of information, facts, and results by the AI system.
Safety (trustllm) and Self-Regulation of (goleman) Encourages safe, healthy conversations, avoiding harm or triggers while supporting user well-being.
Fairness (trustllm), Fairness of (trustindoctor) and Empathy of (goleman) Impartiality and equity, considering diverse perspectives with a positive, action-oriented tone.
Privacy (trustllm) and Confidentiality domain (trustindoctor) Protects human autonomy, identity, and data dignity through secure practices.
Empathy Fidelity domain (trustindoctor), (goleman) Open, honest expression of sympathy for negative situations or approval for positive ones.
Reliability “Reliable item in (client-clinician), Communication domain in (trustindoctor) and Social Skills in (goleman) Responses that build understanding, connection, and offer encouragement or support.
Ethical Guidance (trustllm), Social Skills of (goleman) Ensures AI promotes emotional health responsibly, avoiding manipulation or harm.
Table 5: Trustworthy Dimensions for Emotional Support Tasks
To assess the trustworthiness of AI responses in emotional support, a tailored set of criteria is essential. The Trustworthy Dimensions framework was developed by integrating key insights from four foundational sources: the TrustLLM framework (trustllm), which focuses on LLM trustworthiness; patient-clinician trust studies (client-clinician; trustindoctor); and Goleman’s emotional intelligence principles (goleman), guiding emotional understanding and action. This process involved extracting relevant dimensions, adapting them for emotional sensitivity and ethical integrity. From TrustLLM (trustllm), we adopted Truthfulness, Safety, Fairness, Privacy, and Machine Ethics as technical foundations. Patient-clinician research (client-clinician; trustindoctor) contributed Honesty, Communication, Confidentiality, Fidelity, and Reliability, emphasizing relational ethics. Goleman’s framework (goleman) added Empathy and Social Skills for emotional resonance. These were refined to remove redundancies, yielding seven dimensions: Truthfulness, Safety, Fairness, Privacy, Empathy, Reliability, and Ethical Guidance, balancing technical reliability with human-centered care, as detailed in Table 5.
In the evolving field of AI for emotional support, trustworthiness is vital for user confidence and effective, empathetic assistance. This study constructs a comprehensive Trustworthy Dimensions framework by blending the TrustLLM criteria (trustllm), insights from patient-clinician trust research (client-clinician; trustindoctor), and Goleman’s emotional intelligence principles (goleman), tailored to emotional support needs with a focus on technical reliability, ethical integrity, and emotional sensitivity.
The TrustLLM framework (trustllm) provides a baseline with eight dimensions–Truthfulness, Safety, Fairness, Robustness, Privacy, Machine Ethics, Transparency, and Accountability–for dependable AI. For emotional support, Robustness and Transparency are less critical, while Safety, Fairness, and Privacy gain prominence. Truthfulness ensures emotional authenticity, and Safety fosters psychologically safe dialogues.
Healthcare trust research enriches this foundation. Crits et al. (client-clinician) highlight trust through reliability and truthfulness (e.g., “I trust my doctor/therapist”) and respect (e.g., “I respect my doctor/therapist”) on a 7-point Likert scale, emphasizing relational support. Richmond et al. (trustindoctor) identify Competence, Fidelity, Honesty, Communication, and Confidentiality, with Fidelity prioritizing user interests and Confidentiality protecting disclosures, aligning with emotional support ethics.
Goleman’s framework (goleman) enhances this with Self-Awareness, Self-Regulation, Motivation, Empathy, and Social Skills, where Empathy (understanding user perspectives) and Social Skills (active listening) are key for emotional connection.
The resulting Trustworthy Dimensions integrate these insights. Truthfulness combines Honesty (trustindoctor) and accuracy (trustllm) for credible, congruent responses. Safety (trustllm) focuses on psychological well-being. Fairness (trustindoctor; trustllm) ensures impartiality with a positive tone. Privacy (trustindoctor; trustllm) safeguards autonomy. Empathy (goleman; trustindoctor) reflects user emotions. Reliability (client-clinician; trustindoctor) builds connection. Ethical Guidance (trustllm; goleman) promotes responsible emotional health.
This framework merges technical and human-centric qualities, offering a robust blueprint for trustworthy AI in emotional support. Future studies could validate and refine it, aligning AI with human emotional complexities for enhanced trust and efficacy.
Appendix B Experiments
B.1 Data splits
We use the datasets’ official splits for our experiments. For MESC (smes), we used a fixed train/val/test split at the utterance level: 23,126 / 2,714 / 2,922 for train/val/test, respectively . For DFEW (DFEW), benchmarks follow 5-fold cross-validation on the 12,059 single-labeled video clips (fd1–fd5; in each fold, one-fifth test and the rest train).
B.2 Hyperparameters Settings
In this part, our processes to finetune the LLMs and train the multimodal encoders, projectors, ConvCompreesor and Reinforcement Learning algorithms are introduced. Our experiments are proceeding on 4 $\times$ H100 GPU.
Supervised Fine-tuning (SFT)
With the SFT process, we fine-tuned models using MESC (smes) designed with causal language modeling objective, resulting in a conversational model capable of understanding video, audio, and text data. The fine-tuning process include 25 training epochs with a learning rate of 2e-5. We apply QLoRA (qlora) with a rank of 128 and an alpha of 256, while keeping the pre-trained audio-visual projectors and ConvCom frozen. Moreover, we used Zero3 (zero3) for multi-gpu training and memory optimization. The models was trained with a maximum gradient norm of 1.0, a warm-up ratio of 0.03, and a weight decay of 0.
Multimodal Projectors training
For video understanding, we use SigLIP-So400M-Patch14-384 (DBLP:siglip/iccv/ZhaiM0B23) as the visual encoder. A fixed set of 16 frames is uniformly sampled from each video and encoded, with features passed through a Spatial-Temporal Connector (STC) (videollama2) for effective spatial-temporal representation. For audio, we adopt BEATs (beats), a pretrained model with an acoustic tokenizer. Audio inputs are converted into 128-bin fbank spectrograms to capture rich auditory features. Both visual and audio encoders remain frozen during training, while only their projection layers are fine-tuned to align with the generative model.
ConvCompressor.
For our ConvCompressor experiments, we used the Mamba-370M (mamba) checkpoint as the backbone. Training was performed for 3 epochs on single utterances (learning rate: $2.5\text{e}{-5}$) and 1 epoch on full conversations (learning rate: $1\text{e}{-4}$), using AdamW with a weight decay of 0.01. To stabilize training, we applied a reduce-on-plateau scheduler (factor: 0.5, patience: 1). Due to GPU memory constraints, we adopted 4-bit quantization (nf4) and LoRA (rank: 64, alpha: 128) for parameter-efficient fine-tuning.
Reinforcement Learning Setup
We trained our model using both PPO and GRPO in 16-bit precision (fp16) with a learning rate of $1\text{e}{-5}$. Prior to reinforcement learning, we applied Supervised Fine-Tuning (SFT) to retain task-specific knowledge and ensure stable performance.
B.3 Conversation Compressor Experiments
We conduct a statistical experiment to show the token reduction efficiency of ConvCompressor on the MESC dataset. Specifically, we measure the compression performance across all samples in both training and test splits. Table 6 demonstrates the conversation history compression effectiveness of ConvCompressor on both training and test datasets. The module achieves remarkable token reduction rates of over 98.5% across both splits, significantly reducing the computational burden while preserving essential contextual information. This substantial compression enables efficient processing of extended dialogue histories without compromising the model’s ability to understand conversational context and emotional nuances.
Split Avg. Before Avg. After Reduction %
Train 863.5 12.2 98.59%
Test 872.3 12.1 98.61%
Table 6: ConvCompressor Token Reduction Statistics
Appendix C Evaluation
C.1 Guideline for Human Evaluation
We trained annotators based on the criteria outlined in Table 8, providing sample examples for reference. Additionally, annotators must follow the labeling process detailed in Figure 5. They are required to achieve a sufficiently high inter-annotator agreement during the final test before proceeding to the official labeling phase. The outcomes of this stage will identify the best frameworks based on various criteria.
C.2 Prompt for LLMs Evaluation
We have also designed prompts for LLMs to serve as evaluators based on the criteria outlined in Table A.3. The full prompt is provided in Figure 6.
Appendix D Ablation Study
Setting Task 1 Task 2 Task 3 Task 4
Acc F1 Acc F1 Acc F1 B-2 R-L BS
w/o video 52.0 53.7 90.3 93.2 35.2 30.3 5.4 17.0 85.9
w/o audio 54.2 49.3 94.1 93.2 23.5 21.2 5.9 17.0 86.2
w/o desc 58.1 56.9 99.1 99.00 42.6 40.5 6.1 17.4 86.5
hist. = 0 45.2 40.6 98.9 91.2 32.1 28.3 3.9 13.6 85.2
hist. = 5 50.6 46.8 99.3 99.0 37.2 32.3 5.0 16.0 85.9
hist. = 9 52.2 46.7 99.3 99.0 40.3 38.9 5.6 16.4 86.0
Table 7: Ablation study of MultiMood highlighting the contribution of each modality and historical context. Metrics include BLEU-2 (B-2), ROUGE-L (R-L), and BERTScore (BS).
Impact of Multimodal Information As shown in Table 7, removing video (and by extension, audio) leads to a sharp drop in performance–emotion recognition decreases by 6.1% and system emotion prediction by 25%. Excluding only audio reduces strategy prediction accuracy by 4.8% and slightly increases response perplexity (+0.19), highlighting the critical role of multimodal signals in emotional support tasks.
Effect of Conversation History. Increasing the number of past dialogue turns consistently improves performance. Without history, emotion recognition reaches only 45.16% and BLEU-2 is 3.85. With 5-turn history, these metrics rise to 50.63% and 5.03, respectively. The best results are achieved with full history using the optimized MultiMood (GRPO + SFT) model: 58.60% emotion recognition, 6.18 BLEU-2, and 40.26% strategy accuracy–demonstrating the strong benefit of historical context.
Effect of Video Descriptions. Adding video descriptions (e.g., facial expressions, tone, volume) brings slight gains. The full model with descriptions scores 58.60% on emotion recognition and 39.65 F1 in strategy, with BLEU-2 at 6.18 and BERTScore at 86.80. Without descriptions, metrics dip marginally–suggesting these annotations add nuance but are not critical for overall performance.
Appendix E MultiMood Trustworthy Examples
To evaluate the trustworthiness of our approach, we assessed MultiMood using two randomly selected samples from the MESC dataset (smes). The results, depicted in Figure 7, indicate that MultiMood consistently delivers safer responses, while the fine-tuned VideoLLaMA2 model (videollama2) tends to provide more direct but potentially concerning answers. For the GPT-4 prompt (GPT4), we applied the same methodology to define trustworthiness criteria, including examples of effective and ineffective responses, allowing the model to score responses accordingly.
Example 1: We examined several cases to highlight the varying effectiveness of AI responses in high-stress scenarios. One notable instance involves a patient overwhelmed by intense work pressure, expressing distress with the provocative remark: "If you had to write an evaluation now. If someone held a gun to your head." VideoLLaMA2 response, "Nobody’s forcing me with the gunpoint. Are you the one with the gun?" avoids the patient’s emotional turmoil, adopting a defensive tone with "Nobody," which may signal disengagement and worsen their isolation. In contrast, the MultiMood response, "Sophie, I’m here to help you. Let’s focus on managing your anxiety and finding ways to cope with your emotions without resorting to self-harm," offers a compassionate, supportive approach. By using the patient’s name, reinforcing its supportive role, and suggesting coping strategies, MultiMood effectively de-escalates the situation. However, it could improve by directly addressing the gun imagery, which indicates severe stress or self-harm risk, to ensure the patient feels validated.
Example 2: In the second case, a patient faces a crisis after a disaster, unable to save everyone, and asks, "They’re all gonna die?" VideoLLaMA2 response, "No, not everyone, but a few people," is blunt and honest, potentially deepening the patient’s emotional distress due to its lack of nuance. Conversely, MultiMood response, "I understand that you’re feeling scared about not being able to save everyone and the possibility of losing loved ones. It’s okay to feel afraid, but let’s work together to find ways to cope with these emotions and support each other," empathizes with the patient, adopts their perspective, and uses positive language to encourage them, avoiding mention of losses while promoting mental stability.
Example 3: The third case features a patient in grief after his mom died. VideoLLaMA2 response, "She was sick, Paul. She was too sick for you to save her" states a factual truth, absolving the patient of responsibility, but lacks emotional support, resembling a robotic dismissal of their pain. In contrast, MultiMood response uses gentler phrasing, such as "life is unpredictable," framing it as "part of being human," to help Paul gradually process his grief rather than rushing past it.
Criteria Definition Positive Response Negative Response
Fluency Which response is smoother and simpler? The response should be clear, easy to read, and flows naturally, avoiding complex or technical language to ensure the user feels at ease. “I’m so sorry you’re feeling overwhelmed. Let’s take a moment to breathe together and talk about what’s been going on.” “Your emotional state appears to be suboptimal. Please provide additional information.”
Identification Which bot is better recognizes personal experiences and more relevant response by directly addressing the user’s specific emotions or situation, making the reply feel personalized and meaningful. “It sounds like losing your pet has been really hard. They were a big part of your life.” “Losing a pet is common. Many people experience this.”
Comfort Which response are more reliable, soothing, and supportive? The response should conveys empathy, reassure the user, and fosters a sense of being understood and cared for during emotional challenges. “You’re not alone in feeling this way. I’m here for you, and we can work through this together.” “You should feel better soon. This is a temporary issue.”
Suggestions Which response is more helpfulness and empathy solutions? Respone should offer practical, compassionate, and tailored advice to support the user’s emotional needs and promote coping strategies. “It might help to journal your thoughts or talk to a close friend. Would you like some tips on starting a journal?” “Just try to stay positive and distract yourself with a hobby.”
Overall Which bot excels at providing emotional assistance for navigating life’s tough and upsetting challenges by integrating fluency, identification, comfort, and helpful suggestions into a cohesive, empathetic response. “I hear how tough this is for you, and it’s okay to feel this way. Let’s try a calming exercise together, and I’m here if you want to share more.” “This situation is difficult. You should seek professional help or read about coping strategies online.”
Table 8: Annotation Guidelines for Evaluating Chatbot Responses in Emotional Support
Figure 5: Our human evaluation process.
{promptbox}
You are an advanced evaluator specializing in emotional support AI systems. Your task is to assess and score responses from three different frameworks (Qwen2-7B, Multimood(SFT), and Multimood(SFT+RL)).
Criterias defined in the following trustworthiness dimensions table: {Trustworthiness dimensions table}.
Provide score based on the response’s alignment with the definitions of these dimensions.
Below are the responses to evaluate: {Response}
Provide the scores in a structured format for each framework, e.g.:
Response 1:
  • Truthfulness: [Score]/10
  • Truthfulness: [Score]/10
  • Continue for all dimensions
Ensure your evaluation is consistent, fair, and reflective of the emotional support context.
Figure 6: Prompt used for LLMs evaluation
Figure 7: A comparation between a finetuned videollama2 model and Multimood to evaluate the trustworthiness of the responses.